Chapter 4 Problem 34

A neutron star has a mass of 2.0 × 1030 kg (about the mass of our sun) and a radius of 5.0 × 103 m (about the height of a good-sized mountain). Suppose an object falls from rest near the surface of such a star. How fast would this object be moving after it had fallen a distance of 0.023 m? (Assume that the gravitational force is constant over the distance of the fall and that the star is not rotating.)
Answer #1

Here , for the acceleration due to gravity at star

g = G * M/R^2

g = 6.673 *10^-11 * 2 *10^30/(5 *10^3)^2

g = 5.3384 *10^12 m/s^2

Now, for the speed

speed = sqrt(2 * g * h)

speed = sqrt(2 * 5.3384 *10^12 * 0.023)

speed = 495546 m/s

the speed of the object is 495546 m/s